if e, f, g, and h are consecutive odd integers and e < f < g < h ,then g + h is how much greater than e + f ?
a 2 b 3 c 4 d 5 e 8

Answer by Edwin McCravy(20063)


If e, f, g, and h were 3, 5, 7, and 9, then g+h would be 7+9=16, and e+f would be 3+5=8, and 16 is 8 greater than 8.

If e, f, g, and h were 39, 41, 43, and 45, then g+h would be 43+45=88 and e+f would be 39+41=80, and 88 is 8 greater than 80.

If e, f, g, and h were 77, 79, 81, and 83, then g+h would be 81+83=164 and e+f would be 77+79=156, and 164 is 8 greater than 156.

Any way you do it, the answer is 8, choice e.  You just have to know what
the words "consecutive" and "odd" mean.

Edwin